# PaTAS: A Parallel System for Trust Propagation in Neural Networks Using Subjective Logic

### v2

@@ III-B Subjective Logic Approaches to Trust
Other SL-based methods focus on interpretable trust quantification. A calibration-based approach [33] clusters model outputs into subjective opinions to derive per-prediction trust scores without accessing internal parameters. While simple and model-agnostic, it again assumes trustworthy datasets and neglects input evidence. The DeepTrust framework by Cheng et al. [34] instead adopts a white-box perspective, integrating dataset evidence during training to assess global model trustworthiness. Although holistic, DeepTrust’s use of SL fusion and multiplication operators lacks algebraic consistency, and its formulation of trust backpropagation is specified only at a single-layer level, making its theoretical extension to deeper architectures unclear despite empirical evaluations on complex networks. PaTAS addresses these limitations by introducing a coherent, layer-wise propagation mechanism compatible with deep architectures.
@@ III-C Trust and Uncertainty Propagation
Existing approaches demonstrate growing interest in uncertainty and trust modeling, yet they often neglect the joint influence of input quality, data reliability, and model parameters on prediction trust. The trustworthiness of an output cannot be viewed as a fixed property of the model alone but must depend on the corresponding input and its propagation through the network. PaTAS addresses these gaps by modeling trust as a dynamic property distributed across inputs, parameters, and activations, enabling consistent and interpretable trust propagation that reflects both data quality and network structure.
@@ IV Proposed Method
In our framework, trust reasoning begins at the feature level, where each individual input component is assigned a trust opinion. This design choice provides greater flexibility and fine-grained control, enabling the system to reflect nuanced variations in input reliability across different operational contexts. These initial trust opinions are then injected into the PaTAS, which propagates them through its network. This propagation mechanism ensures that variations in input trust are explicitly carried through to the model’s outputs, enhancing the interpretability and transparency of AI decisions.
IV-A Foundations of Trust-Aware Neural Inference
This section formalizes the structure and behavior of a trust propagation system that mirrors standard neural network computations.
@@ IV-A Foundations of Trust-Aware Neural Inference
  • $f(\cdot)=\{f^{l}(\cdot)\}$ where $f^{l}(\cdot)$ is the activation function for layer $l$ which may vary across each neurons of the layer.
@@
Given a training dataset $\mathcal{D}=\left\{(\mathbf{x}_{i},\mathbf{y}_{i})\right\}_{i=1}^{N}$, used to train the neural network, and a trust assessment function $T(\mathbf{x})$ evaluating the trustworthiness of each features and labels data ($\mathbf{x}_{i}$ and $\mathbf{y}_{i}$), our objective is to compute a corresponding trust opinion on the network output $\mathbf{y\prime}=f_{\Theta}(\mathbf{x})$. To formalize this, we introduce the notion of a parallel trust function.
Definition 4 (Parallel Trust Function).
The parallel trust function of a neural network with parameters $\Theta$, denoted $Pf_{\Theta}$, is a function that mirrors the structure of the network’s feedforward computation $f_{\Theta}$ to propagate trust assessments from input to output. Given a trust evaluation $T(\mathbf{x})$ over an input $\mathbf{x}$, $Pf_{\Theta}(T(\mathbf{x}))$ returns a trust opinion corresponding to the network’s output $\mathbf{y\prime}=f_{\Theta}(\mathbf{x})$. This opinion reflects how trust in the input influences the trust assigned to the prediction, taking into account the architecture and how the parameters of $f_{\Theta}$ were learned.
To effectively construct $Pf_{\Theta}$, we must first understand how the underlying neural network is built.
@@
While the training process optimizes model accuracy, it does not account for how trust in the input data influences trust in the output predictions. Therefore, accuracy alone is not enough to assess trustworthiness of a model. We therefore turn to Subjective Logic as a framework for modeling and propagating trust through neural networks. As a first step, we consider a simple perceptron model to analyze how input trust opinions can be propagated to the output through the network’s structure and parameters. This forms the basis for progressively building the complete formulation of the parallel trust function $Pf_{\Theta}$.
@@ IV-B Perceptron Case: SL Formulation
Assume that an analyst $A$ wants to form a trust opinion $\omega_{y\prime}^{A}$ on the output $y\prime$ of a perceptron. Since $A$ does not directly observe or interact with $y\prime$, their opinion must be inferred indirectly through the output neuron (or output neurons of the network in the general case), denoted $N_{O}$. Specifically, the analyst relies on the trust opinion $\omega_{y\prime}^{N_{O}}$ formed by the output neuron, and holds a referral trust $\omega_{N_{O}}^{A}$, which expresses how much $A$ trusts $N_{O}$ (or the process used by $N_{O}$) for providing good trust opinion on $y\prime$. Fig. 1 illustrates the corresponding STN.
@@
Let $f_{\Theta}$ be the perceptron inference function, and let $x$ be an input with an associated trust opinion $T_{x}$. Given the output $y\prime=f_{\Theta}(x)$, our goal is to construct $Pf_{\Theta}(T_{x})$, the corresponding trust opinion on $y\prime$.
@@
$y\prime=10\times s+100\times n_{r}\,(\text{eventually plus a bias})$ (8)
@@ IV-B1 Solution to Problem 1 - Trust Propagation from Input to Output
Figure 2: Solution for Problem 1
@@
The model parameters $\theta_{1}=10$ and $\theta_{2}=100$ are fully trusted and input feature trust opinions ($T_{s}$ and $T_{n_{r}}$) are available.
@@
$\omega_{x}^{N_{3}}=fusion(\omega_{x}^{10N_{1}},\omega_{x}^{100N_{2}})$
The choice of the fusion operator (that we will later denote by $\oplus$ ) depends on the semantics of $s$ and $n_{r}$ (see Definition 1). In this example, since $s$ and $p$ are not dependent evidence, the fusion operator to use is cumulative fusion.
@@
To complete the solution, we need to go from $\omega_{x}^{N_{3}}$ to $\omega_{y\prime}^{N_{3}}$ ($y\prime$ is the output of the neural network).
One can observe that:
$y\prime=f(x)$
is a deterministic relation between $y\prime$ and $x$. Hence :
@@
$Pf_{\Theta}(T_{x})=Pf_{\Theta}(\begin{pmatrix}T_{s}&T_{n_{r}}\end{pmatrix})=T_{s}\oplus T_{n_{r}}$
@@ IV-B2 Solution for Problem 2 - Impact of Parameter Input Trust on Output Trust
Trust Opinions $T_{\theta_{1}}$ and $T_{\theta_{2}}$ are assigned to the parameters.
@@
Figure 3: Solution for Problem 2
@@
where $\otimes$ is a trust discount operator. This results is consistent with the solution for problem 1 as for fully trusted $T_{\theta}=(1,0,0)$, we have $\omega_{x}^{\theta N}=T_{\theta}\otimes\omega_{x}^{N}=\omega_{x}^{N}$ (for any neuron $N$).
@@
$Pf_{\Theta}(T_{x})=Pf_{\Theta}(\begin{pmatrix}T_{s}&T_{n_{r}}\end{pmatrix})=(T_{\theta_{1}}\otimes T_{s})\oplus(T_{\theta_{2}}\otimes T_{n_{r}})$ (14)
@@ IV-C Trust Nodes and Trust Functions
In order to formalize the construction of $Pf_{\Theta}$, we introduce two fundamental concepts: the Trust Node (Fig. 4(b)) and the Trust Function (Fig. 4(d)). These components provide the basic mechanisms for propagating trust through the structure of a neural network.
@@
  • $T_{f}^{(l)}$ is the trust-equivalent of the activation function. Usually set to identity function.
Collectively, trust nodes form a Trust Nodes Network (TNN), which mirrors the neural network’s architecture. We formalize this in the following Definition.
Definition 7 (Trust Nodes Network).
@@ V PaTAS For Neural Networks
When neural network parameters such as $\theta_{1}$ and $\theta_{2}$ are learned (e.g., by backpropagation), their trustworthiness depends on the data used for training. If the dataset contains mislabeled or biased samples, parameter trust will be compromised. Yet this is only part of the problem. As discussed in Section II-A, Subjective Logic represents trust as a subjective binomial opinion with the three components trust, distrust, and uncertainty. This decomposition allows us to distinguish between different causes of unreliability: distrust may arise from systematic issues such as mislabeled or poisoned data, while uncertainty reflects variability or noise in the data. A central challenge, therefore, is how to exploit this advantage of subjective logic in order to make this distinction in practice.
To address the above challenges, we introduce the Parallel Trust Assessment System (PaTAS), a framework designed to systematically propagate trust assessments through a neural network. PaTAS operates in parallel with the standard neural architecture and maintains a corresponding structure that mirrors the network’s topology. It enables the computation of trust in the output by integrating two key sources:
@@
    the trust in the parameters, derived from the training dataset using a trust assessment function. This includes both the trust in the input features of the training samples and the trust in the corresponding labels.
V-A PaTAS Overview
The PaTAS is designed to continuously evaluate the trustworthiness and preservation of properties, such as accuracy, during the inference of a neural network. For example, an input $x$ might be accurate, unbiased and trustworthy (i.e., high trust score value $T_{x}$), while the corresponding output $y\prime$ may still be unreliable due to biased or poorly calibrated parameters $\theta$ (i.e., low trust score value $T_{y\prime}$). Rather than altering the neural network itself, PaTAS operates alongside it to evaluate the trustworthiness of computations (see Fig. 5). It mirrors the computations through a parallel set of trust nodes, each aligned with a corresponding neuron. These trust nodes use subjective logic reasoning operations to propagate trust values in a way that reflects the neural computations. The detailed operational flow of PaTAS across feedforward, backpropagation, and inference is illustrated in Fig. 6.
Figure 5: Overview of PaTAS Usage
@@ V-A PaTAS Overview
Feedforward (Training Phase): The NN interface extracts trust scores from input features for all data being processed. These scores (F.1) are passed to the Trust Feedforward module, which propagates them through the Trust Nodes Network in parallel with the neural computations. During training, Trust Feedforward also stores the trust scores of all intermediate computations in the NN, to be reused in backpropagation.
Backpropagation (Training Phase): Once the NN computes gradients, the trust scores of the labels (B.1) and the gradients are provided to the Parameter-Trust Update (B.2). Gradients indicate how dataset labels affect parameter updates, while label trust determines whether these updates should be considered reliable. The stored intermediate trust values are also incorporated in this process. Finally, the Parameter-Trust Update module refines the trust values of the Trust Nodes Network parameters (B.3) , aligning them with the evolving learning dynamics.
@@
PaTAS is composed of four main modules: Trust Feedforward, Output-Trust Aggregation, GenIPTA, and Parameter-Trust Update. Among these, the Parameter-Trust Update requires a more detailed treatment, since its design parallels the role of backpropagation in neural network training and involves an elaborate reasoning operations. We therefore dedicate a separate subsection to it.
V-A1 Trust Feedforward
@@
The Trust Feedforward function propagates trust values through the Trust Nodes Network in alignment with the neural network’s inference flow. It mirrors the layer-wise computations of the original model, but operates entirely on trust values. At each layer, trust in the inputs and parameters is combined using the trust operations defined in the Trust Function (including discounting and fusion), capturing how evidence flows through the network. This process produces a trust opinion for each output of the neural network while also storing intermediate trust scores, which are later used by the Parameter-Trust Update.
V-A2 Output-Trust Aggregation

### v3

@@ III-B Subjective Logic Approaches to Trust
Other SL-based methods focus on interpretable trust quantification. A calibration-based approach [37] clusters model outputs into subjective opinions to derive per-prediction trust scores without accessing internal parameters. While simple and model-agnostic, it again assumes trustworthy datasets and neglects input evidence. The DeepTrust framework by Cheng et al. [38] instead adopts a white-box perspective, integrating dataset evidence during training to assess global model trustworthiness. Although holistic, DeepTrust’s use of SL fusion and multiplication operators lacks algebraic consistency. In particular, it maps neural-network addition to Subjective Logic fusion and neural-network multiplication to SL opinion multiplication, even though these two SL operators belong to different algebraic domains, since fusion operates over agents’ opinions whereas multiplication operates over variables, without clear explanation. Moreover, its formulation of trust backpropagation is specified only at a single-layer level, making its theoretical extension to deeper architectures unclear despite empirical evaluations on complex networks.
PaTAS addresses these limitations by introducing a coherent and well motivated, layer-wise propagation mechanism compatible with deep architectures.
@@ III-C Trust and Uncertainty Propagation
Existing approaches demonstrate growing interest in uncertainty and trust modeling to assess trust in neural networks, yet they often neglect the joint influence of input quality, data reliability, and model parameters on prediction trust. The trustworthiness of an output cannot be viewed as a fixed property of the model alone but must depend on the corresponding input and its propagation through the network. PaTAS addresses these gaps by modeling trust as a dynamic property distributed across inputs, parameters, and activations, enabling consistent and interpretable trust propagation that reflects both data quality and network structure.
@@ IV Proposed Method
In our framework, trust reasoning begins at the feature level, where each individual input component is assigned a trust opinion. This design choice provides greater flexibility and fine-grained control, enabling the framework to reflect nuanced variations in input reliability across different operational contexts. For example, certain regions or even individual pixels of an input image could be individually assigned different trust opinions, derived from measurable indicators such as noise levels, blur estimates, or confidence measures produced by the imaging pipeline. These initial trust opinions are then injected into the PaTAS, which propagates them through its network. This propagation mechanism ensures that variations in input trust are explicitly carried through to the model’s outputs, enhancing the interpretability and transparency of AI decisions.
IV-A Foundations of Trust-Aware Neural Inference for PaTAS
This section formalizes the structure and behavior of a trust propagation framework that mirrors standard neural network computations.
@@ IV-A Foundations of Trust-Aware Neural Inference for PaTAS
  • $f(\cdot)=\{f^{l}(\cdot)\}$ where $f^{l}(\cdot)$ is the activation function for layer $l$ which may vary across each neuron of the layer.
@@
Given a training dataset $\mathcal{D}=\left\{(\mathbf{x}_{i},\mathbf{y}_{i})\right\}_{i=1}^{N}$, used to train the neural network, and a trust assessment function $T(.)$ evaluating the trustworthiness of each features and labels data ($\mathbf{x}_{i}$ and $\mathbf{y}_{i}$), our objective is to compute a corresponding trust opinion on the network output $\mathbf{y\prime}=f_{\Theta}(\mathbf{x})$. To formalize this, we introduce the notion of a Parallel Function.
Definition 4 (Parallel Function).
The Parallel Function of a Neural Network with parameters $\Theta$, denoted $\mathit{Pf}_{\Theta}$, is a function that mirrors the structure of the network’s feedforward computation $f_{\Theta}$ to propagate trust assessments from input to output. Given a trust evaluation $T(\mathbf{x})$ over an input $\mathbf{x}$, $\mathit{Pf}_{\Theta}(T(\mathbf{x}))$ returns a trust opinion on the network’s output $\mathbf{y\prime}=f_{\Theta}(\mathbf{x})$. This opinion reflects the trust assigned to the prediction based on trust in the input, and also taking into account the architecture and how the parameters of $f_{\Theta}$ were learned.
To effectively construct $\mathit{Pf}_{\Theta}$, we must first understand how the underlying neural network is built.
@@
While the training process optimizes model accuracy, it does not account for how trust in the input data influences trust in the output predictions. Therefore, accuracy alone is not enough to assess trustworthiness of a model. We therefore turn to Subjective Logic as a formal calculus for modeling and propagating trust through neural networks. As a first step in constructing the parallel function $\mathit{Pf}_{\Theta}$, we consider a simple perceptron model to analyze how input trust opinions can be propagated to the output through the network’s structure and parameters. This forms the basis for progressively building the complete formulation of $\mathit{Pf}_{\Theta}$.
@@ IV-B Perceptron Case: SL Formulation
Assume that an observer $A$ wants to form a trust opinion $\omega_{y\prime}^{A}$ on the output $y\prime$ of a perceptron. Since $A$ does not directly observe or interact with the internal process used to produce $y\prime$, their opinion must be inferred indirectly through the output neuron (or output neurons of the network in the general case), denoted $N_{O}$. Specifically, the observer relies on the trust opinion $\omega_{y\prime}^{N_{O}}$ formed by the output neuron, and holds a referral trust $\omega_{N_{O}}^{A}$, which expresses how much $A$ trusts $N_{O}$ (or how much $A$ trusts the process used by $N_{O}$) for providing good trust opinion on $y\prime$. Fig. 1 illustrates the corresponding STN.
@@
Let $f_{\Theta}$ be the perceptron inference function, and let $x$ be an input with an associated trust opinion $T_{x}$ [^1] We use $T_{x}$ notation instead of $\omega_{x}$ to emphasize that, as input to the framework, only the opinion on the variable is required and no specific agent needs to be represented.. Given the output $y\prime=f_{\Theta}(x)$, our goal is to construct $\mathit{Pf}_{\Theta}(T_{x})$, the corresponding trust opinion on $y\prime$.
@@
$y\prime=10\times s+100\times n_{r}\,$ (8)
@@ IV-B1 Solution to Problem 1 - Trust Propagation from Input to Output
Figure 2: Trust-propagation based solely on input-feature trust
@@
For now, we assume that the model parameters $\theta_{1}=10$ and $\theta_{2}=100$ are fully trusted and input feature trust opinions ($T_{s}$ and $T_{n_{r}}$) are available.
@@
$\omega_{x}^{N_{3}}=\mathit{fusion}(\omega_{x}^{10N_{1}},\omega_{x}^{100N_{2}})$
The choice of the fusion operator (that we will later denote by $\oplus$ ) depends on the semantics of $s$ and $n_{r}$ (see Definition 1). In this example, since $s$ and $p$ represent independent evidence, the fusion operator to use is cumulative fusion.
@@
Since the output value is deterministically related to the input via
$y^{\prime}=f(x),$
the trust opinion on $y^{\prime}$ is the same as the trust opinion already computed for $x$. Thus,
@@
$\mathit{Pf}_{\Theta}(T_{x})=\mathit{Pf}_{\Theta}(\begin{pmatrix}T_{s}&T_{n_{r}}\end{pmatrix})=T_{s}\oplus T_{n_{r}}$
@@ IV-B2 Solution for Problem 2 - Impact of Parameter Input Trust on Output Trust
Trust Opinions $T_{\theta_{1}}$ and $T_{\theta_{2}}$ are assigned to the parameters. The approach used to calculate these parameter-trust values will be discussed in Section V.
@@
Figure 3: Trust-propagation Trust Network including parameter trust.
@@
where $\otimes$ is a trust discounting operator. This results is consistent with the solution for problem 1 as for fully trusted $T_{\theta}=(1,0,0)$, we have $\omega_{x}^{\theta N}=T_{\theta}\otimes\omega_{x}^{N}=\omega_{x}^{N}$ (for any neuron $N$).
@@
$\mathit{Pf}_{\Theta}(T_{x})=\mathit{Pf}_{\Theta}(\begin{pmatrix}T_{s}&T_{n_{r}}\end{pmatrix})=(T_{\theta_{1}}\otimes T_{s})\oplus(T_{\theta_{2}}\otimes T_{n_{r}})$ (14)
@@ IV-C Trust Nodes and Trust Functions
We now need to extend our discussion of $\mathit{Pf}_{\Theta}$ from a single perceptron to larger neural networks. In order to formalize the construction of $\mathit{Pf}_{\Theta}$, we introduce two fundamental concepts: the Trust Node (Fig. 4(b)) and the Trust Function (Fig. 4(d)). These components provide the basic mechanisms for propagating trust through the structure of a neural network.
@@
  • $T_{f}^{(l)}$ is the trust-equivalent of the activation function. In this work, we set it to identity function as we use ReLU as activation function.
@@ V PaTAS For Neural Networks
When neural network parameters such as $\theta_{1}$ and $\theta_{2}$ are learned (e.g., by backpropagation), their trustworthiness depends on the data used for training. If the dataset contains mislabeled or biased samples, parameter trust will be compromised. Yet this is only part of the problem. As discussed in Section II-A, Subjective Logic represents trust as a subjective binomial opinion with the three components trust (belief), distrust (disbelief), and uncertainty. This decomposition allows us to distinguish between different causes of unreliability: distrust may arise from systematic issues such as mislabeled or poisoned data, while uncertainty reflects variability or noise in the data. A central challenge, therefore, is how to exploit this advantage of subjective logic in order to make this distinction in practice.
V-A PaTAS General Description
To address the previous challenges, we introduce the Parallel Trust Assessment System (PaTAS), a framework designed to systematically propagate trust assessments through a neural network. PaTAS operates in parallel with the standard neural architecture and maintains a corresponding structure that mirrors the network’s topology. It enables the computation of trust in the output by integrating two key sources:
@@ V-A PaTAS General Description
    the trust in the parameters, derived from the training dataset trustworthiness calculated using a trust assessment function. This includes both the trust in the input features of the training samples and the trust in the corresponding labels.
To extend the trust-propagation principles introduced for the perceptron in the previous section, PaTAS organizes its Trust Nodes into a structure that mirrors the full neural network. Each neuron is associated with a corresponding Trust Node, and these nodes are connected according to the network’s topology. This generalization forms what we refer to as a Trust Nodes Network (TNN).
Definition 7 (Trust Nodes Network).
The PaTAS is designed to continuously evaluate the trustworthiness and preservation of properties, such as accuracy, during the inference of a neural network. For example, an input $x$ might be accurate, unbiased, and trustworthy (i.e., high trust score value $T_{x}$), while the corresponding output $y\prime$ may still be unreliable due to biased or poorly calibrated parameters $\theta$ (i.e., low trust score value $T_{y\prime}$). Rather than altering the neural network itself, PaTAS operates alongside it to evaluate the trustworthiness of computations (see Fig. 5). Embedding trust operations inside the network would modify its internal computations, risk harming accuracy, and substantially increase the computational cost. For this reason, PaTAS performs trust assessment in parallel, ensuring that trust reasoning does not interfere with the model’s predictions or complicate its operation.
Figure 5: Overview of PaTAS Workflow
V-B PaTAS Design
Figure 6: High-Level Overview of PaTAS Integration with a Neural Network
@@ V-B PaTAS Design
Feedforward (Training Phase): The NN interface extracts trust scores from input features for all data being processed. These trust scores (TF.1) are passed to the Trust Feedforward module, which propagates them through the Trust Nodes Network in parallel with the neural computations. During training, Trust Feedforward also stores the trust scores of all intermediate computations in the NN, to be reused in backpropagation.
Backpropagation (Training Phase): Once the NN computes gradients, the trust scores of the labels (B.1) and the gradients are provided to the Parameter-Trust Update. Gradients (B.2) indicate how dataset labels affect parameter updates, while label trust determines whether these updates should be considered reliable. The stored intermediate trust values are also incorporated in this process. Finally, the Parameter-Trust Update module refines the trust values of the Trust Nodes Network parameters (B.3), aligning them with the evolving learning dynamics.
@@
As illustrated in Fig. 6, PaTAS is composed of four main modules: Trust Feedforward, Output-Trust Aggregation, GenIPTA, and Parameter-Trust Update. Among these, the Parameter-Trust Update requires a more detailed treatment, since its design parallels the role of backpropagation in neural network training and involves elaborate reasoning operations. We therefore dedicate a separate subsection to it.
V-B1 Trust Feedforward
@@
The Trust Feedforward function propagates trust values through the Trust Nodes Network in alignment with the neural network’s inference flow. It mirrors the layer-wise computations of the original model, but operates entirely on trust values. At each layer, trust in the inputs and parameters is combined using the trust operations defined in the Trust Function (including discounting and fusion), capturing how evidence flows through the network. This process produces a trust opinion for each output neuron of the neural network while also storing intermediate trust scores, which are later used by the Parameter-Trust Update.
V-B2 Output-Trust Aggregation